OpenStudy (kitkat16):

There are 6 girls and 7 boys in a class. A team of 10 players is to be selected from the class. If the selection is random, what is the probability of selecting a team of 5 girls and 5 boys?

OpenStudy (kitkat16):

@eliassaab can u help please

OpenStudy (anonymous):

Hold on

OpenStudy (kitkat16):

how did you come up with 286

OpenStudy (anonymous):

286 is the number of all possibilities from your problem 1

OpenStudy (kitkat16):

so what is the equation to fiqure this out?

OpenStudy (anonymous):

To choose k elements form n elements, you compute \[ \frac{n!}{k!(n-k)!} \]

OpenStudy (anonymous):

there are 286 ways to choose 10 out of 13.

OpenStudy (kitkat16):

13/10

OpenStudy (anonymous):

Then you have to choose 5 girls from 7 and 5 boys from 6

OpenStudy (anonymous):

What is is 13/10?

OpenStudy (kitkat16):

10 out of 13

OpenStudy (anonymous):

If you want to choose 2 elements from {a,b,c}, you have {a,b} <{a,c} and {b,c} three ways. \[ \frac {3!}{2! (3-2)!}=3 \]

OpenStudy (kitkat16):

11/25 0.44